jaynator495 (jaynator495):

will medal due to its a difficultish random qustion lol just because i wanted to post a quastion lol 64/256+45x97/5-5-4563*98/72

jaynator495 (jaynator495):

@yayanieves

jaynator495 (jaynator495):

lol

OpenStudy (anonymous):

:/ Lmao idkkkkk

jaynator495 (jaynator495):

i didnt think you would just felt like calling you lol

OpenStudy (mokeira):

\[\frac{ 64 }{256 }+45\times \frac{ 97 }{5 }-5-4563\times \frac{ 98 }{72 }\] is this what you mean

jaynator495 (jaynator495):

i meant 64/256 as division when you see the / i meant division

OpenStudy (mokeira):

oooh ok

jaynator495 (jaynator495):

im guessing his could take a while... lol

OpenStudy (mokeira):

BODMAS \[64\div256+45\times97\div5-5-4563\times98\div72\] start with the division \[0.25+45\times19.4-5-4563\times1.361111111\] next, multiplication \[0.25+873-5-6210.749999\] next addition \[873.25-5-6210.749999\] finish it up \[=-5342.499999\]

OpenStudy (mokeira):

finally done lol

jaynator495 (jaynator495):

i already iknew the steps just wanted to finally post a second quastion lol

OpenStudy (mokeira):

lol!!!!!!!!!!!!!!!! haha nice one

jaynator495 (jaynator495):

ikr xD

jaynator495 (jaynator495):

and the more accuret answer is -5342.5 XD

jaynator495 (jaynator495):

aha i saw it before it was deleted! XD

Elsa213:

My typing is annoying though e.e